OpenStudy (stormswan):

Can someone please check my work?? I will give a medal :) A circle is centered at (7, 8) and has a radius of 11. Which of the following is the equation for this circle? (x − 7)^2 + (y − 8)^2 = 121 (x − 7)^2 + (y − 8)^2 = 11 (x + 7)^2 + (y + 8)^2 = 121 (x + 7)^2 + (y + 8)^2 = 11

OpenStudy (stormswan):

i think its the second one

OpenStudy (anonymous):

The left part is correct. For the right side you have to square the radius

OpenStudy (stormswan):

so.... it would be the first one then?

OpenStudy (anonymous):

yes

OpenStudy (stormswan):

okay, thanks! can you help me with another one?

OpenStudy (anonymous):

ok

OpenStudy (stormswan):

Quadrilateral OPQR is inscribed in circle N, as shown below. Which of the following could be used to calculate the measure of ∠QRO? Circle N is shown with a quadrilateral OPQR inscribed inside it. Angle O is labeled x plus 16. Angle P is not labeled. Angle Q is labeled 6x minus 4. Angle R is labeled 2x plus 16. m∠QPO + (x + 16)° + (6x − 4)° = 360° m∠QPO = (x + 16)° + (6x − 4)° m∠QPO + (2x + 16)° = 180° m∠QPO = (6x − 4)° + (2x + 16)°

OpenStudy (stormswan):

I'm not sure how to find this one though. do you need the picture of it or are you okay?

OpenStudy (anonymous):

yeah i'm gonna need the pic

OpenStudy (stormswan):

okay hold up

OpenStudy (stormswan):

there you go :)

OpenStudy (anonymous):

ok let me think on it for a second.

OpenStudy (anonymous):

the third one is true. opposite angles in an inscribed quadrilateral are supplementary

OpenStudy (stormswan):

that's the answer? how did you get it? can you help me understand? :)

OpenStudy (anonymous):